40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6916     Entry time: Thu Jul 5 01:34:11 2012
Author: yuta 
Type: Update 
Category: Locking 
Subject: MI with X arm ALS 

I tried to lock FPMI using ALS, but I could not take care of ALS for both arms + MI. So, I decided to try one arm + MI.
I don't know why, but I couldn't make it. We need investigation.

Procedure I took:

  1. Align FPMI.

  2. Misalign ETMY.

  3. Press CLEAR HISTORY for C1:ALS-BEATY_FINE_PHASE filter module.
    Are there any command to do this?

  4. Stabilize X arm length.
    I made a script for turning on ALS servo nicely. It currently lives in /users/yuta/scripts/easyALS.py. You have to specify the arm(X or Y) and sign of the gain. It needs to be refined.

  5. Sweep the offset and stabilize X arm lenth to IR resonance.
   (Ran /opt/rtcds/caltech/c1/scripts/ALS/findIRresonance.py Xarm)

  6. Tried to lock MI. I tried to do this by feeding back the signal to BS or ITMs. Both worked fine when ALS holds X arm to IR off-resonance, but I couldn't lock MI when ALS holds X arm to IR resonance. This may come from too much phase fluctuation from X arm reflection. We should investigate this.

Handing off the servo from ALS to LSC:

  I made a script to do this. It just decreases ALS gain and increases LSC gain with 30 sec ramp time. It needs to be refined, so it currently lives in /users/yuta/scripts/handofftoLSC.py. It worked fine without loosing IR transmission.

ALS stability:
  Current stabiliy of the ALS servo is not enough. It doesn't stay for more than ~ 10min. I suspect this is from frequency servo of end lasers losing lock, or beat signals being too small for the beat box because of intensity fluctuation of green transmission. We definitely need to align end greens, but it is painful.

ELOG V3.1.3-